# bibliography.bib

### v3

@@ 1.4 History of $\mathrm{ITM}$’s, $\mathrm{IET}$’s and the dynamics of piecewise isometries
$\mathrm{ITM}$’s are a special class of piecewise isometries. There is a vast literature on such systems, and so we mention just a small subset of it: [MR1905204], [MR1938473], [MR1772421], [MR1738947], [MR2091702], [MR1992662], [MR2039048], [MR2221800], [MR2486783], [MR3010377], [MR4032960], [MR4075314], [MR4441154].
In [MR4082258] the authors connected 1D and 2D piecewise isometries, by embedding $\mathrm{IET}$’s into planar piecewise isometries.
@@ 1.5 Organization of the paper
The structure of the paper is as follows. In Section 2 we prove that $\overline{X}$ is equal to the non-wandering set of and a few other auxiliary results used throughout the paper. In Section 3 we introduce the product notation, define eventually periodic maps and prove Theorem $A$. In Section 4 we prove the crucial Theorem 4.6 about the linear dependence of critical itinerary vectors. In Section 5 we define the ACC and Matching conditions and prove Theorem $B$, using a special case of Theorem 4.6, while in Section 6 we prove Theorem $C$, using the full version of Theorem 4.6. In Section 7 we discuss the parameter space and possible future research. We discuss the connections of ITMs to billiards in Appendix A. Finally, in Appendix B we give discuss and give the link to the GitHub repository of the program written by Björn Winckler which allows one to experiment with the family of ITMs introduced in [MR2013352].
@@ 2.2 The non-wandering set of $T$
2.3 Properties of finite type $\mathrm{ITM}$’s
@@ 3.1 Product notation and perturbations
Let $W(r):=\mathbb{R}{}^{r}\bigoplus\mathbb{R}{}^{r-1}$ be the $(2r-1)$-dimensional real vector space that contains the coefficient vectors (introduced below), and let $(\bm{e}_{1},\dots,\bm{e}_{r})$ and $(\bm{f}_{1},\dots,\bm{f}_{r-1})$ be the canonical bases for $\mathbb{R}{}^{r}$ and $\mathbb{R}{}^{r-1}$, respectively. Recall that $\mathrm{ITM}(r)$ is the parameter space of $\mathrm{ITM}$’s on $r$ intervals, and that it is a convex polytope contained in $\mathbb{R}^{2r-1}$. We call the elements of this space parameter vectors. These vectors also have canonical coordinates coming from the ambient space $\mathbb{R}^{2r-1}$. We will use the shorthand $(\gamma\,\beta)$ for a parameter vector $(\gamma_{1}\dots\gamma_{r}\,\beta_{1}\dots\beta_{r-1})$.
@@
Product notation can also be used to define parameter families of $\mathrm{ITM}$’s. Assume that the parameter vector $(\gamma\,\beta)$ of $T$ satisfies some equation:
@@
Recall that the set $IMT(r)$ of all $\mathrm{ITM}$’s on $r$ intervals can be viewed as a closed convex polytope in $\mathbb{R}^{2r-1}$. Therefore, a family $\mathcal{T}$ given by a set of vectors $\mathcal{E}$ can be viewed as a convex polytope of co-dimension $t=\text{dim}(\text{span}(\mathcal{E}))$ as the intersection of $\mathrm{ITM}(r)$ and the linear sub-space $E$ spanned by $\mathcal{E}$. If $\mathcal{T}$ is now a rational family, then there exists an invertible matrix $A$ with rational coefficients that maps $E$ to a coordinate sub-space, say, $x_{1}=\ldots=x_{2r-1-t}=0$ (where we denoted by $x_{i}$ some Cartesian coordinates in $\mathbb{R}^{2r-1}$). It is clear that the points with rational coordinates are dense in the image $A(\mathcal{T})$. Since the matrix $A$ has rational coefficients, so does $A^{-1}$, and thus the points with rational coordinates are dense in $A^{-1}\cdot A(\mathcal{T})=\mathcal{T}$. Hence, parameters with rational coefficients are dense in the rational family $\mathcal{T}$. ∎
@@ 5.1 Definitions and examples
Theorem 5.12 (Stability is equivalent to ACC + Matching).
@@ 6.1 Unstable number and correspondence
Figure 9: The result of applying $\delta$ to particular $J_{0}$ for which the return map has $5$ branches.

### v4

@@ 1.5 Overview of $\mathrm{IET}$s, $\mathrm{ITM}$s and piecewise isometries
$\mathrm{ITM}$s are a special class of piecewise isometries. There is a vast literature on such systems, and so we mention just a small subset of it: [MR1905204], [MR1938473], [MR1772421], [MR1738947], [MR2091702], [MR1992662], [MR2039048], [MR2221800], [MR2486783], [MR3010377], [MR4032960], [MR4075314], [MR4441154]. In [MR4082258], the authors connected 1D and 2D piecewise isometries, by embedding $\mathrm{IET}$s into planar piecewise isometries.
@@ 1.6 Organization of the paper
The structure of the paper is as follows. In Section 2 we prove that $\overline{X}$ is equal to the non-wandering set of and a few other auxiliary results used throughout the paper. In Section 3 we introduce the product notation, define eventually periodic maps, and prove Theorem $A$. In Section 4 we prove the crucial Theorem 4.6 about the linear dependence of critical itinerary vectors. In Section 5 we define the ACC and Matching conditions and prove Theorem $B$, using a special case of Theorem 4.6, while in Section 6 we prove Theorem $C$, using the full version of Theorem 4.6. In Section 7 we show that the set of stable maps has full measure in the set of finite type maps, which strengthens the Boshernitzan–Kornfeld Conjecture to the Irrational Rotations Conjecture. In Section 8 we discuss the parameter space and possible future research. The connection between $\mathrm{ITM}$s and billiards is contained Appendix A. Finally, in Appendix B we give a link to the GitHub repository of the program written by Björn Winckler which allows one to experiment with the family of $\mathrm{ITM}$s introduced in [MR2013352].
@@ 2.2 The non-wandering set of $T$
2.3 Properties of finite type $\mathrm{ITM}$s
@@ 3.1 Product notation and perturbations
Let $W(r):=\mathbb{R}{}^{r}\bigoplus\mathbb{R}{}^{r-1}$ be the $(2r-1)$-dimensional real vector space that contains the coefficient vectors (introduced below), and let $(\bm{e}_{1},\dots,\bm{e}_{r})$ and $(\bm{f}_{1},\dots,\bm{f}_{r-1})$ be the canonical bases for $\mathbb{R}{}^{r}$ and $\mathbb{R}{}^{r-1}$, respectively. Recall that $\mathrm{ITM}(r)$ is the parameter space of $\mathrm{ITM}$s on $r$ intervals, and that it is a convex polytope contained in $\mathbb{R}^{2r-1}$. We call the elements of this space parameter vectors. These vectors also have canonical coordinates coming from the ambient space $\mathbb{R}^{2r-1}$. We will use the shorthand $(\gamma\,\beta)$ for a parameter vector $(\gamma_{1}\dots\gamma_{r}\,\beta_{1}\dots\beta_{r-1})$.
@@
Product notation can also be used to define parameter families of $\mathrm{ITM}$s. Assume that the parameter vector $(\gamma\,\beta)$ of $T$ satisfies some equation:
@@
Recall that the set $IMT(r)$ of all $\mathrm{ITM}$s on $r$ intervals can be viewed as a closed convex polytope in $\mathbb{R}^{2r-1}$. Therefore, a family $\mathcal{T}$ given by a set of vectors $\mathcal{E}$ can be viewed as a convex polytope of co-dimension $t=\text{dim}(\text{span}(\mathcal{E}))$ as the intersection of $\mathrm{ITM}(r)$ and the linear sub-space $E$ spanned by $\mathcal{E}$. If $\mathcal{T}$ is now a rational family, then there exists an invertible matrix $A$ with rational coefficients that maps $E$ to a coordinate sub-space, say, $x_{1}=\ldots=x_{2r-1-t}=0$ (where we denoted by $x_{i}$ some Cartesian coordinates in $\mathbb{R}^{2r-1}$). It is clear that the points with rational coordinates are dense in the image $A(\mathcal{T})$. Since the matrix $A$ has rational coefficients, so does $A^{-1}$, and thus the points with rational coordinates are dense in $A^{-1}\cdot A(\mathcal{T})=\mathcal{T}$. Hence, parameters with rational coefficients are dense in the rational family $\mathcal{T}$. ∎
@@ 5.1 Definitions and examples
Theorem 5.12 (Characterization of Stability).
@@ 6.1 Unstable number and correspondence
Figure 9: The result of applying $\delta$ to the particular $J_{0}$ for which the return map has $5$ branches.
@@
7 Almost every finite type map is stable
In this section, we show that the Boshernitzan–Kornfeld Conjecture implies the Irrational Rotations Conjecture. In fact, this results follows immediately the Characterization of Stability Theorem 5.12 and the following theorem:
Theorem 7.1.
The set of all infinite type maps has full measure in $ITM(r)\setminus S(r)$.
Proof of Boshernitzan–Kornfeld Conjecture $\implies$ Irrational Rotations Conjecture.
Since infinite type maps have zero measure in $\mathrm{ITM}(r)$, by Theorem 7.1 the stable maps have full measure in $\mathrm{ITM}(r)$. By Theorem 5.12, we know that stable maps correspond to maps for which the return map to every interval is either a rotation or the identity. The set of stable maps with rationally independent coefficients forms a full measure subset of stable maps, and for these maps, every return map corresponds to an irrational circle rotation. Thus the maps corresponding to irrational circle rotations form a full measure subset of $\mathrm{ITM}(r)$. ∎
We now turn to the proof of Theorem 7.1. Until now, we only needed to consider first return maps on intervals contained in $X$, for which it is known that they are bijective and have finitely many branches, and thus correspond to $\mathrm{IET}$s. It is clear that this notion can be generalized to an arbitrary interval $J\subset X$.
Definition 7.2.
Let $T$ be an $\mathrm{ITM}$ and assume that there is an interval $J\subset I$ such that each point of $I$ returns to $J$ under iterates of $T$. Then $J$ is partitioned into maximal intervals on which the itinerary up to the return time to $J$ is constant. The map $R_{J}$ defined piecewise on these intervals as the iterate of $T$ for which this interval returns to $J$ is called the return map to $J$.
Compared to the definition of the return map to a component interval of $X$, the generalized return map does not need to be bijective. Thus the images of its branches are not necessarily disjoint. The next lemma tells us when a generalized return map has a continuation for all sufficiently small perturbations.
Lemma 7.3.
Let $T$ be an $\mathrm{ITM}$ and let $J\subset I$ be an interval such that the generalized return map $R_{J}$ is well-defined. Assume that:
  • For each point $a$ in the interior of $J$, we have that the orbit of $a$ contains at most one critical point of $T$;
  • The boundary points $x,y$ of $J$ do not land on discontinuities for all time.
Then for any sufficiently small perturbation of $T$, there exists an interval $\tilde{J}$ close to $J$ such that the generalized return map $R_{\tilde{J}}$ to $\tilde{J}$ is well-defined.
Proof.
Let $J_{1},\dots,J_{n}$ be the component intervals of $R_{J}$, let $r_{1},\dots,r_{n}$ be their return times to $J$, and let $a_{1},\dots,a_{n-1}$ be the points in the interior of $J$ that land on discontinuities. By the second assumption, we know that $J_{1}$ does not land back on itself when it returns to $J$. Indeed, this would mean that $x$ is a periodic point. Since the points to the left of $x$ are not contained in $X$, this means that $x$ must be the left boundary point of a maximal periodic interval, so it must eventually land on discontinuity, contradicting the second assumption. By an analogous argument, we know that $J_{n}$ also does not land on itself when it returns to $J$, so both $J_{1}$ and $J_{n}$ land into the interior of $J$ when they return to $J$. This means that there is an interval $\hat{J}$ containing $J$, which is larger by a definite size, and such that a generalized return map $R_{\hat{J}}$ on $\hat{J}$ is well-defined. By the first assumption, for every sufficiently small perturbation of $T$, the points $a_{1},\dots,a_{n-1}$ have continuations $\tilde{a}_{1},\dots,\tilde{a}_{n-1}$ which land on the same discontinuities as before and have the same landing times. This means that every interval $\tilde{J}_{i}=[\tilde{a}_{i-1},\tilde{a}_{i})$ has the same itineraries up to time $r_{i}$. Let $\tau$ be the permutation associated to $J$, so that the intervals $J_{\tau(1)}$ and $J_{\tau(n)}$ get mapped to the front and back of $J$, respectively. For a sufficiently small perturbation $\tilde{T}^{r_{\tau(1)}}(\tilde{J}_{\tau(1)})$ and $\tilde{T}^{r_{\tau(n)}}(\tilde{J}_{\tau(n)})$ are both contained in $\hat{J}$. Thus the interval $[\tilde{T}^{r_{\tau(1)}}(a^{+}_{\tau(1)-1}),\tilde{T}^{r_{\tau(n)}}(a^{-}_{\tau(n)}))$ has a well-defined generalized return map and is close to $J$. ∎
We now show a lemma about the global stability of the critical set and critical orbits in $[T_{0}]$. Recall that $\mathcal{C}_{X}(\tilde{T})$ denotes the set of all critical points of $\tilde{T}$ contained in $X(\tilde{T})$. For a $\tilde{\beta}\in\mathcal{C}_{X}(\tilde{T})$, let $J(\tilde{\beta})$ be the component interval of $X(\tilde{T})$ containing $\tilde{\beta}$. Let $r(\tilde{\beta}^{-})$ and $r(\tilde{\beta}^{+})$ be the return times of $\tilde{\beta}^{-}$ and $\tilde{\beta}^{+}$ to $J(\tilde{\beta})$, respectively.
Lemma 7.4.
Let $[T_{0}]$ be the stable region. Then:
  • For any $\tilde{T}\in[T_{0}]$ the set $\mathcal{C}_{X}(\tilde{T})$ is equal to the set of continuations of the discontinuities contained in $\mathcal{C}_{X}(T_{0})$.
  • If the $\tilde{\beta}$ is the continuation of a discontinuity $\beta\in\mathcal{C}_{X}(T_{0})$, then $r(\tilde{\beta}^{-})=r(\beta^{-})$ and $r(\tilde{\beta}^{+})=r(\beta^{+})$.
  • The $\tilde{T}$-itineraries of $\tilde{\beta}^{-}$ and $\beta^{+}$ are equal to the $T_{0}$-itineraries of $\beta^{-}$ and $\beta^{+}$ up to and including the time $r(\beta^{-})+r(\beta^{+})$.
More briefly, we will say that $\mathcal{C}_{X}(T_{0})$, $r(\beta^{-}),r(\beta^{+})$ and the itineraries of $\beta^{+},\beta^{-}$ up to and including the time $r(\beta^{-})+r(\beta^{+})$ are constant in $[T_{0}]$. Note that for a stable map $\tilde{T}$, all of these are locally constant in a small neighbourhood of $\tilde{T}$ in parameter space. The proof is in the first part of the proof of Theorem 5.23. Lemma 7.4 tells us that this is true globally inside the stable region as well.
Proof.
Since $[T_{0}]$ is connected and open, it is path-connected. Let $\eta:[0,1]\to[T_{0}]$ be a path starting at $T_{0}$ and ending at $\tilde{T}$. Since $\eta(t)$ is a stable map, for every $t\in[0,1]$, there is a small neighbourhood $\mathcal{U}(t)$ of $\eta(t)$ such that the critical set, return times and itineraries of critical points up to the times described above are constant. By compactness, the cover of $\eta([0,1])$ by the neighbourhoods $\mathcal{U}(t)$ has a finite subcover $\mathcal{U}(t_{1}),\dots,\mathcal{U}(t_{p})$, where $t_{1}<\dots<t_{p}$. Since these neighbourhoods are open, $\mathcal{U}(t_{i})\cap\mathcal{U}(t_{i+1})$ is non-empty, for all $1\leq i<p$. Thus the critical set, return times, and itineraries are constant on the union of these sets as well. By induction, it must be constant on the entire path $\eta([0,1])$, and thus constant on the entire stable region $[T_{0}]$. ∎
The crucial reason why Theorem 7.1 holds is the following theorem:
Theorem 7.5.
Let $T$ be a finite map in the boundary of stable region. Then $T$ has a critical connection.
This means that the parameters defining such a map must satisfy a linear equation with integer coefficients. Thus such maps live on a countable union of codimension-$1$ subspaces and are therefore of zero measure.
Proof.
Let $[T_{0}]$ be the stable region such that $T\in\partial[T_{0}]$. Let $\{T_{n}\}_{n\in\mathbb{N}}$ be a sequence of maps in $[T_{0}]$ converging to $T$. We may assume that $T$ satisfies the ACC property because otherwise $T$ has a critical connection. We will show that this means that the sequence of non-wandering sets $X(T_{n})$ converges in the Hausdorff topology and that the resulting set is a subset of $X(T)$. This immediately follows if we prove that the constancy from Lemma 7.4 extends to $T$ in the following sense: the continuations of discontinuities in $\mathcal{C}_{X}(T_{0})$ are contained in $\mathcal{C}_{X}(T)$ and their return times and itineraries are the same as for $T_{0}$.
For the sake of contradiction, assume that there is a discontinuity $\beta_{T_{0}}\in\mathcal{C}_{X}(T_{0})$ such that its continuation $\beta_{T}$ has a different itinerary up to the return time of $\beta_{T_{0}}$ to its component interval of $X(T_{0})$. We may without loss of generality assume that $\beta_{T}$ is of $+$-type, with the other case being analogous. Let $k$ be the smallest time for which the itineraries are different. This means that $T_{0}^{k}(\beta_{T_{0}})\in I_{s}$ and $T^{k}(\beta_{T})\in I_{s^{\prime}}$, where by continuity $s^{\prime}\in\{s-1,s+1\}$. Since $\beta_{T}$ is of $+$-type, we must in fact have that $s^{\prime}=s+1$. Indeed, by ACC, none of the images $T_{n}^{k}(\beta_{T_{n}})$ land on $\beta_{s}^{+}(T_{n})$, where $\beta_{T_{n}}$ are the continuations of $\beta_{T_{0}}$. In particular, they are all to the right of $\beta_{s}^{+}$. Thus passing to the limit, we have that $T^{n}(\beta_{T})$ can’t be to the left of $\beta_{s}(T)$, so $s^{\prime}=s+1$. By a similar argument as before, we have that all of the iterates $(\beta_{T_{n}})$ must be strictly to the left of $\beta_{s+1}^{-}(T_{n})$. Thus if the itinerary changes, in the limit we must have that $T_{n}^{k}(\beta_{T_{n}})$ lands on $\beta_{s+1}^{+}(T)$. This is a contradiction with our assumption that there are no critical connections, so $\beta_{T}$ has the same itinerary up to the return time of $\beta_{T_{0}}$ to its component interval of $X(T_{0})$.
A similar argument then shows that each iterates $T^{r(\beta_{T}^{+})}$ has the same itinerary up to time $r(\beta_{T}^{-})$ as $\beta_{T}^{-}$. Indeed, again by looking at the smallest time when the itinerary changes, we get a critical connection for $T$. This clearly shows that the required constancy extends to $T$ in a continuous way, meaning that the orbits of intervals $J_{\beta}$, where $J_{\beta}$ is the interval containing $\beta\in\mathcal{C}_{X}(T_{0})$, move continuously from $[T_{0}]$ to $T$. Finally, we will show that the orbits of different intervals $J_{\beta}$ must remain separated, i.e. they do not touch. Indeed, if two components of different orbits touched, we would have a critical connection. This is because these different components must have different itineraries. Thus they must at some point land on different sides of some discontinuity $\beta_{s}$. By considering the first such time, we must have that the point at which these components touch lands on $\beta_{s}$, thus creating a critical connection, since the boundary points of these components are iterates of critical points. Thus $X(T_{0})$ has a homeomorphic continuation inside $X(T)$ with the same dynamics as $X(T_{0})$.
Since $T$ is not stable, there exists a component interval $J$ of $X(T)$ for which the return map violates Matching, and thus has at least three branches. By our previous discussion, the orbit of $J$ has a definite distance from the continuation of $X(T_{0})$. Indeed, it is clear that is disjoint from it and does not touch the the continuation of $X(T_{0})$ because the touching point would have to land on discontinuity to separate the orbits of the touching components of $X(T_{0})$ and $O(J)$, since they must have different itineraries. Since the ACC condition holds for $T$, the interval $J$ satisfies all of the assumptions of Lemma 7.3. Thus by Lemma 7.3, for any sufficiently small perturbation $\tilde{T}$ of $T$ there is an interval $\tilde{J}$ close to $J$ which has a well-defined generalized return map $R_{\tilde{J}}$. Since $T$ is in the boundary of $J$, we may assume that $\tilde{T}\in[T_{0}]$. But then $\tilde{J}$ contain points in $\tilde{X}$ which are definite distance from the continuation of $X(T_{0})$. This is clearly a contradiction, so $T$ does not satisfy ACC and must therefore have a critical connection. ∎
As a corollary, we get a sufficient condition for a finite type map to not be in the closure of any stable region:
Corollary 7.6.
Let $T$ be a finite type map that does not satisfy Matching and has no critical connections. Then $T$ is not in the closure of any stable region.
Proof.
Since $T$ does not satisfy Matching, it is not contained in any stable region. Since it does not have any critical connection, by Theorem 7.5 it is also not contained in the boundary in any stable region. Thus it is not contained in the closure of any stable region. ∎
Theorem 7.5 allows us to establish the following criterion for a map to be of infinite type:
Lemma 7.7.
Let $T$ be a map that is not a finite type stable map and assume that the defining parameters $(\gamma_{1},\dots,\gamma_{r},\beta_{1},\dots,\beta_{r-1})$ are rationally independent. Then $T$ is an infinite type map.
To prove this criterion, we need the following lemma about the number of equations a map needs to satisfy in order for a return map to an interval of $X$ to have $n\geq 2$ branches.
Lemma 7.8.
Let $T$ be a map such that there is an interval $J$ contained in $X$ on for which the return map $R_{J}$ has exactly $n\geq 2$ continuity intervals. Then there are at least $n-2$ linearly independent vectors $v_{1},\dots,v_{n-2}$ with integer coefficients such that defining parameters $(\gamma\,\beta)$ of $T$ satisfy $\langle v_{i},(\gamma\,\beta)\rangle=0$, for all $21\leq i\leq n-$ 2.
Proof.
Let $\sigma$ be the permutation associated to $R_{J}$ and let $\tau$ be its inverse. We have the following string of $n-1$ equalities that must hold:
$R_{J}(a^{-}_{\tau(1)})$ $=R_{J}(a^{+}_{\tau(2)-1})$
$R_{J}(a^{-}_{\tau(2)})$ $=R_{J}(a^{+}_{\tau(3)-1})$
$\dots$
$R_{J}(a^{-}_{\tau(n-1)})$ $=R_{J}(a^{+}_{\tau(n)-1}),$
where the $a$’s are the discontinuities of $R_{J}$ and the boundary points of $J$. There are two cases. The first is when the boundary points of $J$ land on discontinuities. Then we have that $R_{J}(a^{\pm}_{i})=\langle R^{J,\pm}_{i},(\gamma\,\beta)\rangle$, where $R^{J,\pm}_{i}$ is the corresponding return vector, so all the equalities above are equivalent to:
$\begin{split}\langle R^{J,-}_{\tau(1)}-R^{J,+}_{\tau(2)-1},(\gamma\,\beta)\rangle&=0\\ \langle R^{J,-}_{\tau(2)}-R^{J,+}_{\tau(3)-1},(\gamma\,\beta)\rangle&=0\\ &\dots\\ \langle R^{J,-}_{\tau(n-1)}-R^{J,+}_{\tau(n)-1},(\gamma\,\beta)\rangle&=0\end{split}$ (21)
If there is a linear dependence between the $n-1$ vectors $R^{J,-}_{\tau(1)}-R^{J,+}_{\tau(2)-1},\dots,R^{J,-}_{\tau(n-1)}-R^{J,+}_{\tau(n)-1}$ with coefficients $\alpha_{1},\dots,\alpha_{n-1}$, we can extend it to the linear dependence of all vectors associated to the return map $J$, by setting the coefficients for all other vectors to 00. Then it easily follows from Theorem 4.6 that all of the coefficients of this linear dependence must be zero, so $\alpha_{1}=\dots=\alpha_{n-1}=0$. This is due to the simple fact that return vector $R^{J,+}_{\tau(1)-1}$ does not appear among the vectors in (21), so the coefficient next to it is 00, and this propagates in the conclusion of Theorem 4.6 to all coefficients being equal to zero. Thus in this case we get $n-1$ linearly independent vectors.
If either $a_{0}^{+}$ or $a_{n}^{-}$ does not land on a discontinuity, then we must replace in (21) the vector $R_{0}^{J,+}$ by $R_{\tau(1)-1}^{J,+}+R_{0}^{J,+}$, or the vector $R_{n}^{J,-}$ by $R_{\tau(n)}^{J,-}+R_{n}^{J,-}$. This is because in this case $R_{J}(a^{+}_{0})=a_{0}^{+}+\langle R^{J,+}_{0},(\gamma\,\beta)\rangle$ and $a_{0}^{+}=\langle R_{\tau(1)-1}^{J,+},(\gamma\,\beta)\rangle$, and similarly for $a_{n}^{-}$. If either of these vectors lands on a discontinuity, then the same argument as above shows that we get $n-1$ linearly independent vectors, because either $R_{\tau(1)-1}^{J,+}$ or $R_{\tau(n)}^{J,-}$ will have coefficient zero. Thus we may assume that both boundary points do not land on discontinuities. But then we may simply take the subset of $n-2$ vectors that does not include the vector related to the return of $a_{0}^{+}$, and the same argument as before shows that these vectors are linearly independent. ∎
The last part of the proof where we reduce the set from $n-1$ to $n-2$ vectors might appear unnecessary, but it is simple to check that for cases $n=3$ and $n=2$, the set of $n-1$ vectors we get in the last part of the proof is linearly dependent.
Proof of Lemma 7.7.
Since we have assumed that the defining parameters are rationally independent, we know that $T$ has no critical connections so it satisfies ACC. If $T$ did not satisfy Matching, there would have to be a component interval of $X$ with at least three branches of the return map. By Lemma 7.8 this would mean that there is at least one non-zero vector $v$ with integer coefficients such that $\langle v,(\gamma\,\beta)\rangle$ is equal to 00. This contradicts our assumption on rational independence, so $T$ satisfies Matching. Thus if $T$ were of finite type, it would have to be stable, but we have assumed that this is not the case. Thus $T$ has to be of infinite type. ∎
Finally, Lemma 7.7 allows us to prove Theorem 7.1:
Proof of Theorem 7.1.
Let $\mathrm{INF}(r)$ be the set of all infinite type $\mathrm{ITM}$s on $r$ intervals. The set of all rationally independent parameters $\mathcal{R}(r)$ has full measure in $ITM(r)$. By Lemma 7.8, we know that any map associated with such parameters has to be either a stable finite type map or an infinite type map. Thus $S(r)\cup\mathrm{INF}(r)$ has full measure in $ITM(r)$, so $\mathrm{INF}(r)$ has full measure in $ITM(r)\setminus S(r)$. ∎